log8(x)+log8(x+2)=1
show that the function f(x)=2x+3 is one to one, by using f(a)=f(b)implies that a=b

Answer by edjones(8007)

log8(x)+log8(x+2)=1
8^(log[8](x)) * 8^log[8](x+2)=8^1
x(x+2)=8
x^2+2x-8=0
(x+4)(x-2)=0
x=2 Why isn't x=-4 an answer also?
.
Ed